News

On this page, we list the main news/changes that affect the instrument and telescope and point at science results with the instrument

  • 2014-10-25 a preliminary catalogue of secondary standard stars is available for download from the tools web page.
  • 2014-09-08 to 2014-09-13 VST was out of operations for 6 nights to allow ESO staff at Paranal to work on the M2 unit whose correct functioning has been compromised in recent weeks.
  • 2014-08-26 to 2014-08-30 a techimncal intervention was needed in order to remove a loose part inside the
    the cryostat vessel of OmegaCAM. The observing constraints from 2014-07-07 can be now removed and VST operates normally. 
  • 2014-08-16 VST is back to normal service.
  • 2014-08-08 Stating from the night of 2014-08-07/08 VST is out of operation for a technical intervantion aimed at repairing the adaptor rotator and TCCDs. VST is expected to be back on operation on the 15th of August. 
  • 2014-07-07 On the 04th of July during daily inspection of the telescope, an anomalous noise was heard coming from inside the OMEGACAM's cryostat. For precaution VST was kept out of service on the nights of 04-05 and 05-06. Operations are now resumed with contraint to not observe at zenith distances above 60 deg (allowed exception: polar field for calibrations). An inspection of the cryostat will be perfromed on mid of July.    
  • 2014-05-30 to 2014-06-08 VST was out of service (9.5 nights) due to a damage to the co-rotator driver. After technical intervention the VST is back to normal operations. 
  • 2014-04-26 VST is back to normal service.
  • 2014-04-23 VST out of service due to a problem with M2 Leg#1: the actuator 1 reached the Hardware Lower Limit HLL. Technical intervention to solve the problem is ongoing.
  • 2014-04-06 The installation of the baffle is completed. Unfortunately the science data taken on the first night after the baffle installation were affected by reflections due to the inner plug of the M1 baffle modifications. On the 7th a non reflecting adhesive tissue was installed and the problem was solved.
  • 2014-03-27 The M1 was washed. This resulted in an improvement of ~3.8% in the mirror's reflectivity.
  • 2014-02-02 VST was out of service due to abnormal torque distribution on the azimuth axis range. After technical intervention the telescope was back to normal operations on the night of the 7th.
  • 2014-01-08     Change of flat field morphology due to installation of improved baffling on the VST. The new baffling reduces the effect of stray/scattered light in sky flats and from the moon. A final additional piece of baffling is planned to be installed in February, which will again lead to a change in flat field morphology.
  • 2013-12-05     System throughput drop by about 0.6% within a few days early December, due to dust. Stable since then.
  • 2013-11-20     Small bias level changes in chips 73-76 due to a video board exchange. (Chip 73 was showing occasional strong read out stripes).
  • 2013-11-13 to 2013-11-17 Early technical tests of an enhanced image analysis algorithm, which would use aberrations measured on the science array to complement the intra- and extra-focal measurements on the auxiliary CCDs.
  • 2013-07-21 to 2013-07-25 OmegaCAM was out of operation for 4 nights due to a intervention on VST technical CCD, adapter, and ADC. Not all planned work could be completed and another intervention is foreseen in the next months. Meanwhile VST and OmegaCAM are back to normal operations.
  • 2013-05-07 to 2013-05-13 OmegaCAM was out of operations for six nights due to a cooling leak, which is now recovered and the instrument back on sky as of the night of May 13th.
  • 2013-03-29 to 2013-04-01 OmegaCAM was out of operations for four nights due to a failure of the corotator-rotator alignment which required several days of repair work, both on VST and OmegaCAM side. The system is fully back in operations since 2013-04-02.
  • 2013-03-24 to 2013-03-28 Temporary change in sky flat morphologies due to technical tests with improved baffling. The nights March 24/25 and March 25/26 had one set of additional baffling, the nights March 26/27 and March 27/28 had a different kind of additional baffling. Sky flats taken in the twilights of March 24/25 and March 25/26 are to be used only for science taken on those two nights, analogously skyflats taken in the twilights of 26/27 and March 27/28 are to be used only for science taken in these two nights. As of March 28/29, the baffling was removed and sky flats can be used again to calibrate data taken also before March 24/25. The same applies to dome-flats.
  • 2012-12-05     System throughput increased by 5% due to washing of M1. For science data taken as of December 5/6, 2012, please use flat and standard calibrations taken as of Dec. 6.
  • 2012-11-25     A software update was installed that gets rid of the previously occuring interference noise patterns on science images.
  • 2012-09-12     The BIAS level of Chip 96 was increased from 380 to 650 ADU, to avoid that negative crosstalk from very bright stars in Chip 95 'inversely saturates' the detector at 0 ADU.
  • 2012-09-06     See note in the ESO Call for Proposals for Period 91, Sect. 6.14: filler proposals asking for relaxed observing conditions are especially encouraged (Full Moon, Thin Clouds, Seeing). Possible science cases include proper motion studies, photometric monitoring in single fields, and generally accumulating photons in pre-calibrated fields.
  • 2012-09-01     About 2 nights technical time used by INAF for porting the VST Telescope Control System to the most up-to-date VLT Software version.
  • 2012-07-05     The Illumination Correction Report of the OmegaCAM consortium is available at the bottom of the Manuals section of this webpage.
  • 2012-06-07     Update on Chip 82. On June 2, 2012, the video board connected to Chip 82 was exchanged. Preliminary results show that this has fixed the gain variations. Furthermore, the average gain in terms of ADU/e- has increased by 5 percent (i.e. five percent more flux in flat fields). The video board settings were adopted to not affect the gains of chips 81,83,84 which are controlled by the same board. Note to users : for calibrating chip 82 data taken as of night June 2/3, use only calibrations taken as of June 3 onwards. To be on the safe side, this should also apply for chips 81,83,84, although as mentioned above their gain has remained constant, to within better than 1 percent.
  • 2012-05-09     Note on Chip 82. As mentioned in the OmegaCAM user manual, CCD#82 shows day-to-day gain variations of a few percent, since start of operations. Please be aware of this instability, which occasionally manifests itself in the form of pronounced horizontal step patterns.
  • 2012-03-22 to 2012-03-31 There was an increased occurence of image quality degradation, likely due to a focal plane tilt. About 15 hours of science observations have had to be repeated. The situation has improved after changing M2 positioning parameters, but the root cause is still under investigation.
  • 2012-02-03 to 2012-02-08 Six nights downtime due to OmegaCAM vacuum problem. No significant changes of detector performance have been spotted within the two weeks after recovery. Nevertheless, it is advised that calibrations from after Feb. 8 are not applied blindly to data from before Feb. 3. Please contact usd-help@eso.org in case of doubts.
  • 2012-01-27 to 2012-02-03 A total of 2 nights technical time to work with consortium on OmegaCAM Software action items remaining from PAC.
  • 2012-01-17 The measured seeing distribution on OmegaCAM images from August to December 2011 is linked under the Tools section.
  • 2012-01-06 to 2012-01-10 Successful VST intervention by Paranal engineering for repair and maintenance of rotator brakes, encoders and cooling system.
  • 2011-12-22    Updated version of User Manual and Template Manual for P89 Phase II. See Manuals Page
  • 2011-12-15    Press Release on Science Verification imaging of the spiral galaxy NGC 253.
  • 2011-12-05 to 2011-12-12 VST intervention by INAF to work on remaining Action Items after PAC.
  • 2011-11-05 Preliminary Acceptance Chile is officially granted to VST, retroactively for October 15.
  • 2011-10-15 Official Start of Science Operations, P88
  • 2011-09-14: Updated versions of User Manual and Template Manual, to be used for P88 Phase 2 and for P89 Phase 1. See Manuals Page
  • 2011-08-29 to 2011-09-04: Science Verification, targetting NGC 253 in u+g+r+i and Halpha (NB659).
  • 2011-08-13: Start of Public Surveys with Early Science and Dry Runs in parallel to technical activities
  • 2011-07-28 to 2011-08-12: OmegaCAM commissioning 3
  • 2011-07-13: First Version of Template Manual available from the Manuals Page. Name of User Manual document changed to be in line with ESO standards (no change in contents).
  • 2011-06-29: First version of User Manual available from the Manuals Page.
  • 2011-06-08: Press Release 'First Images from the VLT Survey Telescope'
  • 2011-06-06 to 2011-06-22: OmegaCAM commissioning 2
  • 2011-04-29 to 2011-05-09: OmegaCAM commissioning 1b
  • 2011-03-21 to 2011-03-31: OmegaCAM commissioning 1a